# Strict quantization for compact pseudo-Kähler manifolds and group actions

### v1

@@ 1 Introduction
$\Omega^{0,q}_{k}(X):=\left\{u\in\Omega^{0,q}(X);\,Ru=ik\,u\right\},\ \ q=0,1,2,\ldots,n,$
and
$\Omega^{0,q}_{k}(X)^{G}=\left\{u\in\Omega^{0,q}(X)^{G};\,Ru=ik\,u\right\},\ \ q=0,1,2,\ldots,n.$
@@
$S^{(q)}_{G,k}:L^{2}_{(0,q)}(X)\rightarrow({\rm Ker\,}\Box^{q}_{b})^{G}_{k}$
@@
$H^{G}(X):=({\rm Ker\,}\Box^{n_{-}}_{b})^{G}$
and
$H^{G}_{k}(X):=({\rm Ker\,}\Box^{n_{-}}_{b})^{G}_{k}\,.$
Eventually, for every $G$ and $S^{1}$ invariant function $f\in{C}^{\infty}(X)$, the $k$-th $G$-invariant Toeplitz operator is given by
$T^{G}_{k}[f]:=S^{(n_{-})}_{G,k}\circ f\cdot\circ S^{(n_{-})}_{G,k}:L^{2}_{(0,n_{-})}(X)\rightarrow H^{G}_{k}(X)\,.$ (1)
@@
The contact moment map associated to the form $\omega_{0}$ is the map $\mu:X\to\mathfrak{g}^{*}$ such that, for all $x\in X$ and $\xi\in\mathfrak{g}$, we have
$\langle\mu(x),\xi\rangle=\omega_{0}(\xi_{X}(x))\,.$ (2)
Let $b$ be the nondegenerate bilinear form on $HX$ such that
$b(\cdot,\cdot)=\mathrm{d}\omega_{0}(\cdot,J\cdot)\,.$ (3)
Let us denote $\mu^{-1}(0)$ by $Y$. Assume that 00 is a regular value of $\mu$, assume that the action of $G$ on $\mu^{-1}(0)$ is free and
[…]
where
$\underline{\mathfrak{g}}={\rm Span\,}(\xi_{X};\,\xi\in\mathfrak{g})$
and
$\underline{\mathfrak{g}}^{\perp_{b}}=\left\{v\in HX;\,b(\xi_{X},v)=0,\ \ \forall\xi_{X}\in\underline{\mathfrak{g}}\right\}\,.$
Then, $\mu^{-1}(0)$ is a $d$-codimensional submanifold of $X$. Let $Y:=\mu^{-1}(0)$ and let $HY:=HX\cap TY$. Note that if the Levi form is positive at $Y$, then (4) holds.
Remark 1.2.
Under notation as in Remark 1.1, the action of $G$ on $X$ defined above descends to a Hamiltonian and holomorphic action on $(M,\,\omega,\,J_{M})$. For each $\xi\in\mathfrak{g}$, we denote by $\xi_{M}$ the corresponding infinitesimal vector field on $M$. The moment map will be denoted by
$\Phi\,:\,M\rightarrow\mathfrak{g}^{*}\,.$
Vice-versa, we recall that given a pseudo-Kähler manifold $(M,\,\omega,\,J)$, whose $\omega$ defines an integral cohomology class. Here, $X$ is the circle bundle inside the dual of the quantizing line bundle, see Remark 1.1. A Hamiltonian and holomorphic action of a compact Lie group $G$ on $M$ produces an infinitesimal action of its Lie algebra on $X$. Under suitable topological conditions, for example if the group is semisimple or the manifold $M$ is simply connected, this action can be integrated to an action of $G$ on the circle bundle $X$ inside the dual of the line one, $L^{*}$. This action of $G$ on $X$ preserves $\omega_{0}$ and $J$.
Now, let $R$ be the Reeb vector field of the contact form $\omega_{0}$ and let $\mathcal{L}$ denote the Lie derivative. A vector field $V\,:\,X\rightarrow TX$ such that $\mathcal{L}_{V}\omega_{0}=g\,\omega_{0}$ for some smooth function $g\,:\,X\rightarrow\mathbb{R}$ is called a contact vector field.
By [MDS, Lemma 3.5.14], for every function $g\,:\,X\rightarrow\mathbb{R}$ there exists a unique contact vector field $X_{g}$ which satisfies
$\omega_{0}(X_{g})=g$
and
$\mathrm{d}\omega_{0}(X_{g},\,V)=(\mathrm{d}g(R))\,\alpha(V)-\mathrm{d}g(V)$
@@
$\{g,\,h\}:=\omega_{0}([X_{g},\,X_{h}])=\mathrm{d}g(X_{h})-\mathrm{d}h(X_{g})+\mathrm{d}\omega_{0}(X_{g},\,X_{h})\,$
for $f$ and $g$ smooth function on $X$. Now, we are interested in the subspace of $H$-invariant smooth functions on $X$. Thus, for any $S^{1}$-invariant smooth function $g$ we have $\mathrm{d}g(R)=R(g)=0$ and
$\mathrm{d}\omega_{0}(X_{g},\,V)=-\mathrm{d}g(V)\,.$
Eventually, we note that for two $H$-invariant smooth functions $g$ and $h$ on $X$, we obtain
$\{g,\,h\}=-\mathrm{d}\omega_{0}(X_{g},\,X_{h})+\mathrm{d}\omega_{0}(X_{h},\,X_{g})+\mathrm{d}\omega_{0}(X_{g},\,X_{h})=\mathrm{d}\omega_{0}(X_{h},\,X_{g})\,.$
The CR reduction is the CR manifold $X_{G}:=\mu^{-1}(0)/G$, [HH]. Let $\pi_{X}\,:\,\mu^{-1}(0)\rightarrow X_{G}$ be the projection. Then, $X_{G}$ is a circle bundle over the symplectic reduction $M_{G}:=\Phi^{-1}(0)/G$. In Section 4 we explain how to generalize some of the results in [GS] to the setting of pseudo-Riemannian manifolds. In particular there is an identification
$M_{G}\cong M_{s}/G^{\mathbb{C}}$
where $G^{\mathbb{C}}$ is the complexification of the group $G$ and $M_{s}$ is the saturation by $G^{\mathbb{C}}$ of $\Phi^{-1}(0)$ in $M$. Furthermore, we shall extend the theorem “quantization commutes with reduction” to the pseudo-Riemannian setting. This theorem is well-known and has been proved in different contexts; we refer to [Ma] for a survey. We have decided to include this result here because we explain how to generalize some results along the way of the proof, which may be of interest for further study, see Section 4.
@@
$\lVert\widetilde{f}\rVert_{\infty}:=\lVert f\rVert_{\infty}:=\sup\left\{\lvert\widetilde{f}(x)\rvert\,:\,x\in\mu^{-1}(0)\right\}$ (5)
which is the uniform norm of $f$ on the quotient $M_{G}$. Vice-versa, let $f$ be a smooth $S^{1}$ invariant function on $X_{G}$, its lift defines a smooth $G$ and $S^{1}$ invariant function on $Y$. By [Lee, Lemma 5.34 (b)], the extension Lemma for functions on submanifolds, since $Y$ is a smooth properly embedded submanifold of $X$, then there exists a smooth function $\widetilde{f}$ on $X$ whose restriction to $Y$ coincides with $f$.
As the last piece of notation, let $\widetilde{f}$ be a $G$-invariant function on $X$, we denote the operator norm of its corresponding level $k$ $G$-invariant Toeplitz operator:
$\left\lVert T^{G}_{k}[\widetilde{f}]\right\rVert=\sup\left\{\frac{\lVert T^{G}_{k}[\widetilde{f}]s\rVert}{\lVert s\rVert}\,:\,s\in H^{G}_{k}(X)\,,s\neq 0\right\}\,.$ (6)
In the following theorem we state an analogue of the main result in [BMS]. Parts (c) and (b) of the following theorem are a consequence of the results contained in [GH1], see also [MM2] and [MM1] where the Bergman kernels and the corresponding Toeplitz operators on symplectic manifolds, for the mixed curvature case, are studied. Here, we prove Part (a) by adapting the strategy of [S2] to our setting.
@@
    $\lVert{f}\rVert_{\infty}-\frac{C}{k}\leq\left\lVert T^{G}_{k}[\widetilde{f}]\right\rVert\leq\lVert{f}\rVert_{\infty}\,$
    and in particular
    $\lim_{k}\left\lVert T^{G}_{k}[\widetilde{f}]\right\rVert=\lVert{f}\rVert_{\infty}\,;$
@@
    the semi-classical Dirac condition holds:
    $\left\lVert ki\left[T^{G}_{k}[\widetilde{f}],\,T^{G}_{k}[\widetilde{g}]\right]-T^{G}_{k}[\{\widetilde{f},\,\widetilde{g}\}]\right\rVert=O\left(\frac{1}{k}\right)\,;$
@@
    the semi-classical von Neumann’s condition holds:
    $\left\lVert T^{G}_{k}[\widetilde{f}]\,T^{G}_{k}\left[\widetilde{g}\right]-T^{G}_{k}\left[\widetilde{f}\cdot\widetilde{g}\right]\right\rVert=O\left(\frac{1}{k}\right)\,.$
Furthermore, the Berezin-Toeplitz quantization is also a positive quantization in the following sense, see [LeF]. Let $f\in C^{\infty}(X_{G})^{S^{1}}$ be such that $f(x)\geq 0$ for every $x\in X_{G}$. Then $T_{k}^{G}[f]$ is a non-negative operator, in the sense that
$\left(T_{k}^{G}[f]s,\,s\right)\geq 0$
for every $s\in H_{k}^{G}(X)$. Equivalently, by the spectral theorem, a self-adjoint linear operator $A$ is positive if and only if $\mathrm{Spec}(A)\subset[0,\,+\infty)$. On the other hand, the Weyl quantization of a non-negative function need not to be a positive function.
Definition 1.2.
A star product for the algebra $C^{\infty}(X_{G})^{S^{1}}$ is given by the power series
$g\star h=\sum_{j=0}^{+\infty}C_{j}(g,h)\,\nu^{-j}$
such that $\star$ is an associative $\mathbb{C}[[\nu]]$-linear product, that is,
$(g\star h)\star l=g\star(h\star l)\,,$
for all $g,h,l\in C^{\infty}(X_{G})^{S^{1}}$ and
$C_{0}(g,h)=g\cdot h\,,$
$C_{1}(g,h)-C_{1}(g,h)=i\,\{g,h\}\,,$
for all $g,h\in C^{\infty}(X_{G})^{S^{1}}$. The star product is said to be of Wick type if and only if one function is differentiated in holomorphic directions while the other is differentiated in anti-holomorphic directions only.
It is an easy consequence of [GH1] that the star product defined by Toeplitz operators is of Wick type. Our result prove that $\star$ defines a strict deformation quantization in the sense of [L, Chapter 2]. In the following Corollary we state the properties of the star product induced by Theorem 3.1. Here, we introduce a deformation quantization for $S^{1}$-invariant functions on $X_{G}$ induced by $G$ and $S^{1}$-invariant Toeplitz operators on $X$.
Corollary 1.1.
There exists a unique formal star-product
$f\star g:=\sum_{j=0}^{+\infty}\nu^{j}\,C_{j}(f,\,g)\,,\quad C_{j}(f,\,g)\in C^{\infty}(X_{G})^{S^{1}}$
such that for every smooth functions $f$ and $g$ on $X_{G}$ invariant under the circle action, for every $N\in\mathbb{N}$ we have with suitable constants $K_{N}(f,\,g)$, and for all $k\in\mathbb{N}$
$\left\lVert T^{G}_{k}[f]\,T^{G}_{k}[g]-\sum_{j=0}^{N-1}k^{-j}\,T^{G}_{k}[C_{j}(f,g)]\right\rVert=K_{N}(f,\,g)\,k^{-N}\,.$
Furthermore the star product $\star$ has the following properties.
  • (Unit)
    Let $1$ be the constant function equals to $1$ everywhere, then
    $1\star f=f\star 1=f$
    for every circle invariant smooth function on $X_{G}$.
  • (Parity)
    For every circle invariant smooth functions $f$ and $g$ on $X_{G}$, then
    $\overline{f\star g}=\overline{f}\star\overline{g}\,.$
  • (Trace)
    Denote the trace on $H(X)^{G}_{k}$ by $\mathrm{Tr}_{k}$. It induces a $\mathbb{C}[[\nu]]$-linear map
    $\mathrm{Tr}\,:\,{C}^{\infty}(X_{G})^{S^{1}}[[\nu]]\rightarrow\nu^{-n}\,\mathbb{C}[[\nu]]$
    such that
    $\mathrm{Tr}(f\star g)=\mathrm{Tr}(g\star f)\,.$
Now, if we put $G=\{e\}$, where $e$ is the identity element, we recover previous results in [S1], and we generalize it to pesudo-Kähler manifolds.

### v2

@@ 2.3 CR formulation
$\Omega^{0,q}_{k}(X):=\left\{u\in\Omega^{0,q}(X)\mid Ru={\rm i}ku\right\},\qquad q=0,1,2,\dots,n.$
@@
$T_{k}^{X}[f]:=S^{(n_{-})}_{k}\circ f\cdot\circ S^{(n_{-})}_{k}\colon\ L^{2}_{(0,n_{-})}(X)\rightarrow H_{k}(X).$
Now, recall that $X$ is the circle bundle inside the dual of the line bundle $(L,h_{L})$. We have defined two important operators: $T_{k}^{X}[f]$ (constructed with the $k$-th Fourier component of the Szegő projector $S^{(n_{-})}_{k}$) and $T_{k}[f]$ (constructed with the Bergman projector $P_{k}^{(n_{-})}$). They can be identified, in fact we claim
$S_{k}^{(n_{-})}(x_{m},x^{\prime}_{m^{\prime}})=(x_{m})^{k}\left(\frac{1}{2\pi}P_{k}^{(n_{-})}(m,m^{\prime})(x^{\prime}_{m^{\prime}})^{k}\right),$
where $\pi(x_{m})=m$, $\pi(x^{\prime}_{m^{\prime}})=m^{\prime}$.
Proof of the claim.
We adopt the following conventions and notations: every element $x^{\vee}$ in ${(L^{\vee})^{\vee}=L}$ defines a map $x^{*}\colon L^{\vee}\rightarrow\mathbb{C}$ which is the dual of an element $x\in L^{\vee}$. We also write $f_{m}$ for the restriction of $f\in L^{2}_{(0,n_{-}),k}(X)$ along the fibre $S_{m}$ of the standard circle action on $X$. First, notice that any function $f\in L^{2}_{(0,n_{-}),k}(X)$ is uniquely defined at $m$ along the fibre ${S_{m}\subset L^{\vee}}$ by $f\bigl{(}{\rm e}^{-{\rm i}\theta}\cdot x_{m}\bigr{)}={\rm e}^{{\rm i}k\theta}f(x_{m})$; and thus it defines a linear map ${f_{m}\colon(L^{\vee}_{m})^{\otimes k}\rightarrow T^{*0,q}X}$, which is the same as an element $\sigma(m)\in T^{*0,q}_{m}M\otimes L^{\otimes k}_{m}$. We have the following isomorphism:
$\mathfrak{S}_{k}\colon\ L^{2}_{(0,n_{-}),k}(X)\rightarrow L^{2}_{0,n_{-}}\bigl{(}M,L^{\otimes k}\bigr{)},\qquad f\mapsto\sigma_{f}$
where the inverse of $\mathfrak{S}_{k}$ is given by $\sigma\mapsto f_{\sigma}$, with $f_{\sigma}(x_{m})=(x_{m})^{k}(\sigma(m))$. Finally, the operator given by taking the $k$-th Fourier component along the fibers $S$ of $\pi\colon X\rightarrow M$ is given explicitly by
$F_{k}\colon\ L^{2}_{(0,n_{-})}(X)\rightarrow L^{2}_{(0,n_{-}),k}(X),\qquad f\mapsto\frac{1}{2\pi}\int_{S}f(x)(x^{*})^{k}{\rm d}x.$
Now, the claim follows by the formula $S_{k}^{(n_{-})}=\mathfrak{S}^{-1}_{k}\circ P_{k}^{(n_{-})}\circ\mathfrak{S}_{k}\circ F_{k}$. ∎
2.4 Group action framework
A Hamiltonian and holomorphic action of a compact Lie group $G$ on $M$ induces an infinitesimal action of its Lie algebra on $X$ via the Kostant formula, see [17]. Under suitable topological conditions—such as when $G$ is semisimple or the manifold $M$ is simply connected—this infinitesimal action can be integrated to an action of $G$ on the circle bundle $X$ inside the dual line bundle $L^{*}$. This action of $G$ on $X$ preserves both $\omega_{0}$ and $J$. We are thus led to introduce the main definitions in the CR setting, but before doing so we give more details on the Kostant formula and the quantization of group actions on pseudo-Kähler manifolds.
Let $M$ be a pseudo-Kähler manifold with complex structure $J$ and Hodge form $\omega$. Recall that we denote by $X$ the circle bundle lying in the dual line bundle $L^{*}$, with circle action ${{\rm e}^{{\rm i}\theta}\cdot\colon S^{1}\times X\rightarrow X}$, projection $\pi\colon X\rightarrow M$ and $\partial_{\theta}$ the generator of the structure circle action on it. Hence $X$ is naturally a contact and Cauchy–Riemann manifold; $\omega_{0}$ is the contact form.
Suppose given, in addition, an action $\mu\colon G\times M\rightarrow M$ of a compact Lie group $G$ with Lie algebra $\mathfrak{g}$, which is holomorphic with respect to the complex structure $J$ and Hamiltonian with moment map $\Phi\colon M\rightarrow\mathfrak{g}^{\vee}$, $2\omega(\xi_{M},\cdot)={\rm d}\Phi^{\xi}(\cdot)$. By [17], the action $\mu$ naturally induces an infinitesimal contact action of $\mathfrak{g}$ on the circle bundle $X$. Explicitly, if $\xi\in\mathfrak{g}$ and $\xi_{M}$ is the corresponding Hamiltonian vector field on $M$, then its contact lift $\xi_{X}$ is as follows. Let $\mathbf{v}^{\sharp}$ denote the horizontal lift on $X$ of a vector field $\mathbf{v}$ on $M$, we have
$\xi_{X}:=\xi_{M}^{\sharp}-\langle\Phi\circ\pi,\xi\rangle R.$ (2.1)
Furthermore, suppose that the infinitesimal action (2.1) can be integrated to an action of $G$ on $X$. By hypothesis, it preserves the Cauchy–Riemann structure and the contact form $\alpha$. There is a naturally induced unitary representation of $G$ on the Hardy space $H(X)\subset L^{2}_{(0,n_{-})}(X)$ given by $g\colon f\mapsto f\circ{g^{-1}}\cdot$.
Thus, adopting the convection of the previous subsection, we find that $X$ admits an action of a compact connected Lie group $G$ of dimension $d$, and the $G$ action preserves $\omega_{0}$ and the CR structure. Furthermore, for any $\xi\in\mathfrak{g}$, the vector field $\xi_{X}$ appearing in equation (2.1) is $(\xi_{X}u)(x)=\frac{\partial}{\partial t}\bigl{[}u\bigl{(}{\rm e}^{t\xi}x\bigr{)}\bigr{]}_{|_{t=0}}$, for any $u\in{C}^{\infty}(X)$.
@@ 2.4 Group action framework
$g^{*}\colon\ \Lambda^{r}_{x}(T^{*}X\otimes\mathbb{C})\rightarrow\Lambda^{r}_{g^{-1}\circ x}(T^{*}X\otimes\mathbb{C})$
be the pullback map. Since $G$ preserves $J$, we have $g^{*}\colon T^{*0,q}_{x}X\rightarrow T^{*0,q}_{g^{-1}\circ x}X$ for all $x\in X$. Thus, for $u\in\Omega^{0,q}(X)$, we have $g^{*}u\in\Omega^{0,q}(X)$. Put
$\Omega^{0,q}(X)^{G}:=\big{\{}u\in\Omega^{0,q}(X)\mid g^{*}u=u,\,\forall g\in G\big{\}}.$
[…]
$L^{2}_{(0,q)}(X)^{G}:=\big{\{}u\in L^{2}_{(0,q)}(X)\mid g^{*}u=u,\,\forall g\in G\big{\}}.$
Put $\bigl{(}\operatorname{Ker}\Box^{q}_{b}\bigr{)}^{G}:=\operatorname{Ker}\Box^{q}_{b}\cap L^{2}_{(0,q)}(X)^{G}$. The $G$-invariant Szegő projection is the orthogonal projection
$S^{(q)}_{G}\colon\ L^{2}_{(0,q)}(X)\rightarrow\bigl{(}\operatorname{Ker}\Box^{q}_{b}\bigr{)}^{G}$
with respect to (|)($\cdot$|$\cdot$)( ⋅ | ⋅ ). Let $S^{(q)}_{G}(x,y)\in{D}^{\prime}\bigl{(}X\times X,T^{*0,q}X\boxtimes\bigl{(}T^{*0,q}X\bigr{)}^{*}\bigr{)}$ be the distributional kernel of $S^{(q)}_{G}$, ${D}^{\prime}$ denotes the space of distribution sections, see Section 3.2 for notation.
Let $f\in C^{\infty}(X,\mathbb{C})$ be invariant under the circle action and the action of $G$, then the $G$-invariant Toeplitz operator is given by
$T^{G}[f]:=S^{(n_{-})}_{G}\circ f\cdot\circ S^{(n_{-})}_{G}\colon\ L^{2}_{(0,n_{-})}(X)\rightarrow\bigl{(}\operatorname{Ker}\Box^{n_{-}}_{b}\bigr{)}^{G}.$
The transversal circle action can be used to define the corresponding Fourier components. First, we note that ${\rm e}^{{\rm i}\theta}\cdot(gx)=g\bigl{(}{\rm e}^{{\rm i}\theta}\cdot x\bigr{)}$, for every $x\in X$, $\theta\in[0,2\pi]$, $g\in G$. For every $k\in\mathbb{Z}$, let
$\Omega^{0,q}_{k}(X)^{G}=\big{\{}u\in\Omega^{0,q}(X)^{G}\mid Ru={\rm i}ku\big{\}},\qquad q=0,1,2,\dots,n.$
[…]
$S^{(q)}_{G,k}\colon\ L^{2}_{(0,q)}(X)\rightarrow\bigl{(}\operatorname{Ker}\Box^{q}_{b}\bigr{)}^{G}_{k}$
with respect to (|)($\cdot$|$\cdot$)( ⋅ | ⋅ ). For ease of notation, we put $H^{G}(X):=\bigl{(}\operatorname{Ker}\Box^{n_{-}}_{b}\bigr{)}^{G}$ and $H^{G}_{k}(X):=\smash{\bigl{(}\operatorname{Ker}\Box^{n_{-}}_{b}\bigr{)}^{G}_{k}}$. Eventually, for every $G$ and $S^{1}$ invariant function $f\in{C}^{\infty}(X)$, the $k$-th $G$-invariant Toeplitz operator is given by
$T^{G}_{k}[f]:=S^{(n_{-})}_{G,k}\circ f\cdot\circ S^{(n_{-})}_{G,k}\colon\ L^{2}_{(0,n_{-})}(X)\rightarrow H^{G}_{k}(X).$
2.5 A key result on Toeplitz operators
The goal of this subsection is to establish a key result concerning Toeplitz operators. This result will serve as the main analytical tool in the proof of the principal theorem of the paper.
@@ 2.5 A key result on Toeplitz operators
The contact moment map associated to the form $\omega_{0}$ is the map $\mu\colon X\to\mathfrak{g}^{*}$ such that, for all $x\in X$ and $\xi\in\mathfrak{g}$, we have $\langle\mu(x),\xi\rangle=\omega_{0}(\xi_{X}(x))$.
Note that $\mu$ is $\Phi\circ\pi\colon X\rightarrow\mathfrak{g}^{*}$. Let $b$ be the nondegenerate bilinear form on $HX$ such that ${b(\cdot,\cdot)={\rm d}\omega_{0}(\cdot,J\cdot)}$. Clearly $b$ is the pullback via $\pi$ of the pseudo-Riemannian metric on $M$. Let us denote $\mu^{-1}(0)$ by $Y$. Assume that 00 is a regular value of $\mu$, assume that the action of $G$ on $\mu^{-1}(0)$ is free and
$\underline{\mathfrak{g}}_{x}\cap\underline{\mathfrak{g}}^{\perp_{b}}_{x}=\left\{0\right\}\qquad\text{at every point}\ x\in Y,$ (2.2)
where $\underline{\mathfrak{g}}={\rm Span}(\xi_{X};\xi\in\mathfrak{g})$ and $\underline{\mathfrak{g}}^{\perp_{b}}=\left\{v\in HX\mid b(\xi_{X},v)=0,\,\forall\xi_{X}\in\underline{\mathfrak{g}}\right\}$. Note that (2.2) is one of the main assumptions in [11]. By [11, Section 2.5], we observe that when restricted to ${\mathfrak{g}\times\mathfrak{g}}$, the bilinear form $b$ is nondegenerate on $Y$. Note that in Theorem 1.5 we assume that the pseudo-metric $g$ restricted to the orbit through any $m\in\Phi^{-1}(0)$ is non-degenerate. This assumption implies equation (2.2). Then, $\mu^{-1}(0)$ is a $d$-codimensional submanifold of $X$. Let $Y:=\mu^{-1}(0)$ and let $HY:=HX\cap TY$.
Now, recall that $R$ is the Reeb vector field of the contact form $\omega_{0}$ and let $\mathcal{L}$ denote the Lie derivative. A vector field $V\colon X\rightarrow TX$ such that $\mathcal{L}_{V}\omega_{0}=g\omega_{0}$ for some smooth function $g\colon X\rightarrow\mathbb{R}$ is called a contact vector field. By [25, Lemma 3.5.14], for every function $g\colon X\rightarrow\mathbb{R}$ there exists a unique contact vector field $X_{g}$ which satisfies $\omega_{0}(X_{g})=g$ and
${\rm d}\omega_{0}(X_{g},V)=({\rm d}g(R))\alpha(V)-{\rm d}g(V)$
@@
$\{g,h\}:=\omega_{0}([X_{g},X_{h}])={\rm d}g(X_{h})-{\rm d}h(X_{g})+{\rm d}\omega_{0}(X_{g},X_{h})$
for $f$ and $g$ smooth function on $X$. Now, we are interested in the subspace of $H$-invariant smooth functions on $X$. Thus, for any $S^{1}$-invariant smooth function $g$ we have ${\rm d}g(R)=R(g)=0$ and ${{\rm d}\omega_{0}(X_{g},V)=-{\rm d}g(V)}$. Eventually, we note that for two $H$-invariant smooth functions $g$ and $h$ on $X$, we obtain
$\{g,h\}=-{\rm d}\omega_{0}(X_{g},X_{h})+{\rm d}\omega_{0}(X_{h},X_{g})+{\rm d}\omega_{0}(X_{g},X_{h})={\rm d}\omega_{0}(X_{h},X_{g}).$
The CR reduction is the CR manifold $X_{G}:=\mu^{-1}(0)/G$ [11]. Let $\pi_{X}\colon\mu^{-1}(0)\rightarrow X_{G}$ be the projection. Then, $X_{G}$ is a circle bundle over the symplectic reduction $M_{G}:=\Phi^{-1}(0)/G$. In Section 5.1, we explain how to generalize some of the results in [9] to the setting of pseudo-Riemannian manifolds. In particular, there is an identification $M_{G}\cong M_{s}/G^{\mathbb{C}}$, where $G^{\mathbb{C}}$ is the complexification of the group $G$ and $M_{s}$ is the saturation by $G^{\mathbb{C}}$ of $\Phi^{-1}(0)$ in $M$.
@@
$\big{\lVert}\widetilde{f}\big{\rVert}_{\infty}:=\lVert f\rVert_{\infty}:=\sup\big{\{}\big{\lvert}\widetilde{f}(x)\big{\rvert}\mid x\in\mu^{-1}(0)\big{\}},$
which is the uniform norm of $f$ on the quotient $M_{G}$. Vice versa, let $f$ be a smooth $S^{1}$ invariant function on $X_{G}$, its lift defines a smooth $G$ and $S^{1}$ invariant function on $Y$. By [20, Lemma 5.34 (b)], the extension Lemma for functions on submanifolds, since $Y$ is a smooth properly embedded submanifold of $X$, then there exists a smooth function $\widetilde{f}$ on $X$ whose restriction to $Y$ coincides with $f$.
As the last piece of notation, let $\widetilde{f}$ be a $G$-invariant function on $X$, we denote the operator norm of its corresponding level $k$ $G$-invariant Toeplitz operator
$\big{\lVert}T^{G}_{k}\big{[}\widetilde{f}\big{]}\big{\rVert}=\sup\left\{\frac{\lVert T^{G}_{k}\big{[}\widetilde{f}\big{]}s\rVert}{\lVert s\rVert}\mid s\in H^{G}_{k}(X),\,s\neq 0\right\}.$ (2.3)
In the following theorem, we state an analogue of the main result in [1]. Parts (c) and (b) of the following theorem are a consequence of the results contained in [7], see also [23, 24], where the Bergman kernels and the corresponding Toeplitz operators on symplectic manifolds, for the mixed curvature case, are studied. In Section 6, we prove part (a) by adapting the strategy of [33] to our setting.
@@
    $\lVert{f}\rVert_{\infty}-\frac{C}{k}\leq\big{\lVert}T^{G}_{k}\big{[}\widetilde{f}\big{]}\big{\rVert}\leq\lVert{f}\rVert_{\infty}$
    and in particular $\lim_{k}\big{\lVert}T^{G}_{k}\big{[}\widetilde{f}\big{]}\big{\rVert}=\lVert{f}\rVert_{\infty}$;
@@
    the semi-classical Dirac condition holds
    $\big{\lVert}k\mathrm{i}\big{[}T^{G}_{k}\big{[}\widetilde{f}\big{]},T^{G}_{k}\big{[}\widetilde{g}\big{]}\big{]}-T^{G}_{k}\big{[}\big{\{}\widetilde{f},\widetilde{g}\big{\}}\big{]}\big{\rVert}=O\left(\frac{1}{k}\right);$
@@
    the semi-classical von Neumann’s condition holds
    $\big{\lVert}T^{G}_{k}\big{[}\widetilde{f}\big{]}T^{G}_{k}\big{[}\widetilde{g}\big{]}-T^{G}_{k}\big{[}\widetilde{f}\cdot\widetilde{g}\big{]}\big{\rVert}=O\left(\frac{1}{k}\right).$
A main consequence of Theorem 2.3 is Theorem 1.3 stated in the introduction. This is shown in Section 7. Finally, note that the Berezin–Toeplitz quantization is also a positive quantization in the following sense, see [19]. Let $f\in C^{\infty}(X_{G})^{S^{1}}$ be such that $f(x)\geq 0$ for every $x\in X_{G}$. Then $T_{k}^{G}[f]$ is a non-negative operator, in the sense that $\bigl{(}T_{k}^{G}[f]s,s\bigr{)}\geq 0$ for every $s\in H_{k}^{G}(X)$. Equivalently, by the spectral theorem, a self-adjoint linear operator $A$ is positive if and only if ${\mathrm{Spec}(A)\subset[0,+\infty)}$. On the other hand, the Weyl quantization of a non-negative function need not be a non-negative operator.